. Harry's Candy shop’s total monthly revenue increased from $1,000 to $2,000 when its quantity
sold rises from 100 to 300 candies a month. What is the price elasticity of demand for Harry's
candies?
a. 2.318
b. 0.395
c. 2.530
d. 1.816
1
Expert's answer
2019-05-21T08:40:25-0400

Price change from 1000/100=10 to 2000/300=6.67

Price elasticity of demand;

((300-100)/(300+100))/((6.67-10)/(6.67+10))=2.530